Circular No. 6559

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

PSR J0633+1746 A. D. Kuzmin and B. Y. Losovsky, Lebedev Physical Institute, Mos-

cow, report a radio detection of the pulsar PSR J0633+1746 (Geminga) with period 0.237 s. A flux density of ∼ 0.1 Jy at frequency 102 MHz, corresponding to a spectral index of > 1.7, and a dispersion measure of 3 ± 1 pc cm^{−3} was determined. The width of the integrated profile at 50 percent of peak intensity, W_{50}, is 80^{o} of the total pulse phase; W_{10} = 130^{o}. There is apparently an interpulse of 0.2 of the main pulse intensity spaced by ∼ 190^{o}.

GRB 970111 S. R. Kulkarni and M. R. Metzger, California Institute of Technology;

and D. A. Frail, National Radio Astronomy Observatory, report the following in regard to the radio source reported on IAUC 6545 (now denoted as VLA J1528.7+1945) and located in the reduced error box (Hurley et al., IAUC 6545) of GRB 970111: “On Jan. 16 UT, imaging and spectroscopic observations at optical wavelengths of this localization region were obtained using the Keck II 10-m telescope. Two sources, S1 and S2, are identified to lie within 1′′ of VLA J1528.7+1945; their R magnitudes are 20.8 and 21.3, respectively. S1 is a slightly extended galaxy with a strong blue continuum and a redshift of 0.636 (as inferred from the Balmer and Ca II absorption features). S2 is a galaxy located 2′′.15 east and 1′′.06 north of S1. The redshift of S2, based on absorption and [O II] 372.8-nm emission features, is 0.458.”

A. Guarnieri, C. Bartolini, and A. Piccioni, University of Bologna; and N. Masetti, University of Padua, communicate: “Using the Bologna University 1.5-m telescope on Jan. 15.19, 17.19, and 31.15 UT, we observed an object with a mean mag of R = 20.6 ± 0.1 that is practically coincident with the position of the radio source detected by Frail et al. (IAUC 6545). Some indication of variability on a 2-day time scale (to be confirmed) was present. The object is barely visible in our V frame (V − R ∼ 0.0) and is invisible in the B band (B > 21); this would indicate a reddened blue object. On the Palomar plates (epoch 1950.294), the object is undetectable in B and is about at the same R luminosity level stated above. The coincidence of the radio source with this optical object suggests the latter as the probable optical counterpart of both the x-ray and radio sources.”

1997 February 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green